# supReferences

### v4

@@ A.3.1 Proof of Proposition 1
Let $(c_{t})_{t\geq 0}$, $(\rho_{t})_{t\geq 0}$, $(\sigma_{t})_{t\geq 0}$, $C_{\star}\in(0,\infty)$ and $\epsilon_{\star}\in(0,1)$ be as in Lemma 5, and for all $a\in\mathbb{R}$ let $f_{a}=\mathbbm{1}_{(-\infty,a]}$ and, for all $r\in\mathbb{R}$ and $s\in(0,\infty)$, let $h_{a,r,s}$ be as defined in (11) and let $p_{a,s}=\Phi(a/s)$.
@@
By combining this latter result and (55), and by using the fact that $\sup_{x\in\mathbb{R}}|x|G_{q}(x)\leq q^{-1/2}e^{-1/2}$ for all $q\in(0,\infty)$, it follows that
$\big{|}\hat{\eta}_{t+1}(f_{a})-p_{a,\sigma_{t}}\big{|}\leq\big{|}\Psi_{G_{c_{t}}}(\hat{\eta}_{1})(h_{a,\rho_{t},\sigma_{t}})\big{|}\leq\frac{|\rho_{t}|e^{-1/2}}{\hat{\eta}_{1}(G_{c_{t}})\sigma_{1}\sqrt{2\pi c_{t}}}\leq\frac{|\rho_{t}|\sqrt{1+c\sigma^{2}}}{\hat{\eta}_{1}(G_{c_{t}})\sigma_{1}\sqrt{c\rho^{2}}},\quad\forall t\geq 1$ (56)
where the last inequality uses Lemma 5. By Lemma 5 we have $G_{c_{t}}(x)\geq G_{\rho^{2}/\sigma^{2}}(x)$ for all $x\in\mathbb{R}$ and $|\rho^{t}|\leq C_{\star}\epsilon_{\star}^{t}$ for all $t\geq 1$, and thus, by (56),
@@
On the other hand, using the mean value theorem, for all $t\geq 1$ there exists a $v_{t,a}\in([-1,1])$ such that
@@
Noting that in the above calculations $a\in\mathbb{R}$ is arbitrary, by combining (57), (58) and (59) we obtain that
$\|\hat{\eta}_{t+1}-\mathcal{N}_{1}(0,\sigma^{2}_{\infty})\|=\sup_{a\in\mathbb{R}}\big{|}\hat{\eta}_{t+1}(f_{a})-p_{a,\sigma_{\infty}}\big{|}\leq\big{(}C_{1}+C_{\star}/\sigma_{1}\big{)}\epsilon_{\star}^{t},\quad\forall t\geq 1.$
@@
$\|\eta_{t+1}-\mathcal{N}_{1}(0,\rho\sigma^{2}_{\infty}+\sigma^{2})\|\leq\|\eta_{t}-\mathcal{N}_{1}(0,\sigma^{2}_{\infty})\|,\quad\forall t\geq 1.$ (60)
To this aim we let $\hat{\eta}_{\infty}=\mathcal{N}_{1}(0,\sigma^{2}_{\infty})$ and note that, for all $t\geq 2$, we have
$\|\eta_{t}-\mathcal{N}_{1}(0,\rho\sigma^{2}_{\infty}+\sigma^{2})\|$ $=\|\hat{\eta}_{t-1}M-\hat{\eta}_{\infty}M\|=\sup_{a\in\mathbb{R}}|\hat{\eta}_{t-1}(h_{a,\rho,\sigma})-\hat{\eta}_{\infty}(h_{a,\rho,\sigma})|.$
@@
$\sup_{a\in\mathbb{R}}|\hat{\eta}_{t-1}(h_{a,\rho,\sigma})-\hat{\eta}_{\infty}(h_{a,\rho,\sigma})|\leq\|\hat{\eta}_{t-1}-\hat{\eta}_{\infty}\|$
and (60) follows. The proof of the proposition is complete. ∎
A.3.2 Proof of Lemma 3
@@
Let $\delta\in(0,1)$ and remark that, since $\{U_{1}^{n}\}_{n=1}^{N}$ is a set if $N$ independent $\mathcal{U}(0,1)^{s}$ random numbers, it follows from Aistleitner and Hofer, (2014, Theorem 1) that for all $\gamma\in(0,1)$ we have
$\mathbb{P}\Big{(}D^{*}_{s}\big{(}\{U_{1}^{n}\}_{n=1}^{N}\big{)}\leq\delta\Big{)}\geq 1-\gamma,\quad\forall N\geq N_{s,\delta,\gamma}:=\frac{160s+33\log(1/\gamma)}{\delta^{2}}.$
Therefore, since ($(\{U_{t}^{n}\}_{n=1}^{N})_{t\geq 1}$ is assumed to be a sequence of independent sets, it follows that for all $\gamma\in(0,1)$ we have
$\mathbb{P}\Big{(}\sup_{t\in\{1,\dots,T\}}D^{*}_{s}\big{(}\{U_{t}^{n}\}_{n=1}^{N}\big{)}\leq\delta\Big{)}$ $=\mathbb{P}\Big{(}D^{*}_{s}\big{(}\{U_{1}^{n}\}_{n=1}^{N}\big{)}\leq\delta\Big{)}^{T}\geq(1-\gamma)^{T},\quad\forall N\geq N_{s,\delta,\gamma}.$
We now let $q\in(0,1)$ so that, by applying the latter result with $\gamma=1-(1-q)^{1/T}$, we have
$\mathbb{P}\Big{(}\sup_{t\in\{1,\dots,T\}}D^{*}_{s}\big{(}\{U_{t}^{n}\}_{n=1}^{N}\big{)}\leq\delta\Big{)}\geq 1-q,\quad\forall N\geq N_{s,\delta,1-(1-q)^{1/T}}.$
Using Bernoulli’s inequality (see e.g. Li and Yeh,, 2013), we have $1-(1-q)^{1/T}\geq T^{-1}q$ and thus
$N_{s,\delta,1-(1-q)^{1/T}}\leq\frac{160s-33\log(T^{-1}q)}{\delta^{2}}=\frac{160s+33\log(T)+33\log(1/q)}{\delta^{2}}$
and the proof of the lemma is complete. ∎
A.3.3 Proof of Theorem 1
@@
We start with two preliminary results. To state the first one let $\pi_{M}(\mathrm{d}x)=\sum_{m=1}^{M}W_{m}\delta_{\{x_{m}\}}\in\mathcal{P}(\mathbb{R})$ for some $M\in\mathbb{N}$, for all $x\in\{x_{1},\dots,x_{M}\}$ let $\mu_{x}\in\mathcal{P}(\mathbb{R})$, let $\mu=\sum_{m=1}^{M}W_{m}\mu_{x_{m}}\in\mathcal{P}(\mathbb{R})$, $\{u^{n}\}_{n=1}^{N}$ be a point set in $(0,1)$ and $\mu^{N}=N^{-1}\sum_{n=1}^{N}\delta_{F^{-1}_{\mu}(u^{n})}$. Then, it is direct to see that $\|\mu^{N}-\mu\|=D^{*}_{1}(\{u^{n}\}_{n=1}^{N})$. To state the second preliminary result let ($(\{U_{t}^{n}\}_{n=1}^{N})_{t\geq 1}$ be as in the statement of the theorem, and let $\mu_{1}^{N}=\eta_{1}^{N}$ and

### v5

@@ 5.3 Proof of Proposition 1
Let $(c_{t})_{t\geq 0}$, $(\rho_{t})_{t\geq 0}$, $(\sigma_{t})_{t\geq 0}$, $C_{\star}\in(0,\infty)$ and $\epsilon_{\star}\in(0,1)$ be as in Lemma 5, and for all $a\in\mathbb{R}$ let $f_{a}=\mathbbm{1}_{(-\infty,a]}$ and, for all $r\in\mathbb{R}$ and $s\in(0,\infty)$, let $h_{a,r,s}$ be as defined in (28) and $p_{a,s}=\Phi(a/s)$.
@@
By combining this latter result and (36), and by using the fact that $\sup_{x\in\mathbb{R}}|x|G_{z}(x)\leq z^{-1/2}e^{-1/2}$ for all $z\in(0,\infty)$, it follows that, for all $t\geq 1$
$\begin{split}\big|\hat{\eta}_{t+1}(f_{a})-p_{a,\sigma_{t}}\big|\leq\big|\Psi_{G_{c_{t}}}(\hat{\eta}_{1})(h_{a,\rho_{t},\sigma_{t}})\big|\leq\frac{|\rho_{t}|e^{-1/2}}{\hat{\eta}_{1}(G_{c_{t}})\sigma_{1}\sqrt{2\pi c_{t}}}\leq\frac{|\rho_{t}|\sqrt{1+c\sigma^{2}}}{\hat{\eta}_{1}(G_{c_{t}})\sigma_{1}\sqrt{c\rho^{2}}}\end{split}$ (37)
where the last inequality uses Lemma 5. By Lemma 5 we have $G_{c_{t}}(x)\geq G_{\rho^{2}/\sigma^{2}}(x)$ for all $x\in\mathbb{R}$ while $|\rho^{t}|\leq C_{\star}\epsilon_{\star}^{t}$ for all $t\geq 1$, and thus, by (37),
@@
On the other hand, using the mean value theorem, for all $t\geq 1$ there exists a constant $v_{t,a}\in[-1,1]$ such that
@@
Noting that in the above calculations the constant $a\in\mathbb{R}$ is arbitrary, by combining (38), (39) and (40) we obtain that
$\|\hat{\eta}_{t+1}-\mathcal{N}(0,\sigma^{2}_{\infty})\|=\sup_{a\in\mathbb{R}}\big|\hat{\eta}_{t+1}(f_{a})-p_{a,\sigma_{\infty}}\big|\leq\big(C_{1}+C_{\star}/\sigma_{1}\big)\epsilon_{\star}^{t},\quad\forall t\geq 1.$
@@
$\|\eta_{t+1}-\mathcal{N}(0,\rho^{2}\sigma^{2}_{\infty}+\sigma^{2})\|\leq\|\hat{\eta}_{t}-\mathcal{N}(0,\sigma^{2}_{\infty})\|,\quad\forall t\geq 1.$ (41)
To this aim we let $\hat{\eta}_{\infty}=\mathcal{N}(0,\sigma^{2}_{\infty})$ and note that, for all $t\geq 1$, we have
$\|\eta_{t+1}-\mathcal{N}(0,\rho^{2}\sigma^{2}_{\infty}+\sigma^{2})\|$ $=\|\hat{\eta}_{t}M-\hat{\eta}_{\infty}M\|=\sup_{a\in\mathbb{R}}|\hat{\eta}_{t}(h_{a,\rho,\sigma})-\hat{\eta}_{\infty}(h_{a,\rho,\sigma})|.$ (42)
@@
$\sup_{a\in\mathbb{R}}|\hat{\eta}_{t}(h_{a,\rho,\sigma})-\hat{\eta}_{\infty}(h_{a,\rho,\sigma})|\leq\|\hat{\eta}_{t}-\hat{\eta}_{\infty}\|$
and (41) then follows from (42). The proof of the proposition is complete. ∎
5.4 Proof of Lemma 1
5.4.1 Two preliminary results
From Lemma 4 we readily obtain the following discrepancy bound for 1-dimensional importance sampling:
Corollary 3.
@@ 5.4.1 Two preliminary results
The result follows from Lemma 4 and the fact that, for all interval $B=(-\infty,a)\subset\mathbb{R}$, we have
The next lemma is the key intermediate result used to prove Lemma 1:
Lemma 6.
Let $I\subseteq\mathbb{R}$ be a compact set. Then, there exists a constant $C_{I}\in(1,\infty)$ such that, for any sequence $(\mu_{t})_{t\geq 1}$ and letting $\hat{\mu}_{t}=\Psi_{G}(\mu_{t})$ for all $t\geq 1$, for any $T\geq 1$ we have
$\ \max\Big\{\sup_{t\in\{1,\dots,T\}}\|\hat{\mu}_{t}-\hat{\eta}_{t}\|,\sup_{t\in\{1,\dots,T\}}\|\mu_{t}-\eta_{t}\|\Big\}\leq\frac{C_{I}}{\iota_{T-1}}\Delta_{T}\log\big(1+\Delta_{T}^{-1}\big)$
where $\iota_{T-1}=\inf_{t\in\{1,\dots,T-1\}}\mu_{t}(I)$ and $\Delta_{T}=\sup_{t\in\{1,\dots,T\}}\|\mu_{t}-\hat{\mu}_{t-1}M)\|$, with the convention that $\iota_{0}=1$ and that $\hat{\mu}_{0}M=\eta_{1}$.
@@
Let $(\mu_{t})_{t\geq 1}$ and $(\hat{\mu}_{t})_{t\geq 1}$ be as in the statement of the lemma, $(c_{t})_{t\geq 0}$, $(\rho_{t})_{t\geq 0}$ and $(\sigma_{t})_{t\geq 0}$ be as defined in Lemma 5, and for all $a\in\mathbb{R}$ let $f_{a}=\mathbbm{1}_{(-\infty,a]}$ and
Next, we let $\phi:\mathcal{P}(\mathbb{R})\rightarrow\mathcal{P}(\mathbb{R})$ be such that $\phi(\eta)=\Psi_{G}(\eta M)$ for all $\eta\in\mathcal{P}(\mathbb{R})$, and for all $t\geq 1$ we let $\phi^{t}=\phi^{t-1}\phi$ with the convention that $\phi^{0}(\eta)=\eta$ for all $\eta\in\mathcal{P}(\mathbb{R})$. Note that $\hat{\eta}_{t+1}=\phi^{t}(\hat{\eta}_{1})$ for all $t\geq 1$.
Using this notation, and following the approach introduced by Del Moral, (2004, Chapter 7) for studying PFs, we have, using the convention that $\phi(\hat{\mu}_{0})=\hat{\eta}_{1}$,
$\hat{\mu}_{t}-\hat{\eta}_{t}$ $=\sum_{p=1}^{t}\Big(\phi^{t-p}(\hat{\mu}_{p})-\phi^{t-p}\big(\phi(\hat{\mu}_{p-1})\big)\Big),\quad\forall t\geq 1$ (43)
and we now study the terms inside the sums.
To this aim, we assume for now that $T\geq 3$ , and we let $t\in\{2,\dots,T\}$ and $p\in\{1,\dots,t-1\}$ be fixed. Then, for all $a\in\mathbb{R}$ we have, by Lemma 5,
$\begin{split}\phi^{t-p}(\hat{\mu}_{p})(f_{a})&=\frac{\hat{\mu}_{p}Q^{t-p}(f_{a})}{\hat{\mu}_{p}Q^{t-p}(\mathbb{R})}=\Psi_{G_{c+c_{t-p}}}(\mu_{p})(g_{a,t-p})\end{split}$ (44)
and
$\phi^{t-p}\big(\phi(\hat{\mu}_{p-1}))(f_{a})=\Psi_{G_{c+c_{t-p}}}\big(\hat{\mu}_{p-1}M\big)(g_{a,t-p}).$ (45)
By combining (44) and (45), we obtain that
$\begin{split}\sup_{a\in\mathbb{R}}\big|\phi^{t-p}(\hat{\mu}_{p})(f_{a})-&\phi^{t-p}\big(\phi_{p}(\hat{\mu}_{p-1})\big)(f_{a})\big|\\ &=\sup_{a\in\mathbb{R}}\Big|\Psi_{G_{c+c_{t-p}}}(\mu_{p})(g_{a,t-p})-\Psi_{G_{c+c_{t-p}}}\big(\hat{\mu}_{p-1}M\big)(g_{a,t-p})\Big|\\ &=\sup_{a\in\mathbb{R}}\Big|\Psi_{G_{c+c_{t-p}}}(\mu_{p})(h_{a,t-p})-\Psi_{G_{c+c_{t-p}}}\big(\hat{\mu}_{p-1}M\big)(h_{a,t-p})\Big|.\end{split}$ (46)
To proceed further we note that, by Lemma 3, we have $\sup_{(a,k)\in\mathbb{R}\times\mathbb{N}}V(h_{a,k})\leq 1$ and thus, by Lemma 4,
$\begin{split}\sup_{a\in\mathbb{R}}\Big|\Psi_{G_{c+c_{t-p}}}(\mu_{p})(h_{a,t-p})-\Psi_{G_{c+c_{t-p}}}&\big(\hat{\mu}_{p-1}M\big)(h_{a,t-p})\Big|\leq\Big\|\Psi_{G_{c+c_{t-p}}}(\mu_{p})-\Psi_{G_{c+c_{t-p}}}\big(\hat{\mu}_{p-1}M\big)\Big\|.\end{split}$ (47)
Noting that
it follows from Corollary 3 that
$\begin{split}\big\|\Psi_{G_{c+c_{t-p}}}(\mu_{p})-\Psi_{G_{c+c_{t-p}}}\big(\hat{\mu}_{p-1}M\big)\Big\|\leq\frac{5\|\mu_{p}-\hat{\mu}_{p-1}M\|}{\mu_{p}(G_{c+c_{t-p}})}&\leq\frac{5\|\mu_{p}-\hat{\mu}_{p-1}M\|}{\mu_{p}(G_{c+c_{\star}})}\leq\frac{5\Delta_{T}}{\mu_{p}(G_{c+c_{\star}})}\end{split}$ (49)
where the second inequality uses the fact that, by Lemma 5, we have $\sup_{k\geq 1}c_{k}\leq c_{\star}:=\rho^{2}/\sigma^{2}$. Noting that since $p<T$ we have $\mu_{p}(G_{c+c_{\star}})\geq\iota^{\prime}\iota_{T-1}$, with $\iota^{\prime}=\inf_{x\in I}G_{c+c_{\star}}(x)>0$, it follows from (49) that
$\begin{split}\big\|\Psi_{G_{c+c_{t-p}}}(\mu_{p})-\Psi_{G_{c+c_{t-p}}}\big(\hat{\mu}_{p-1}M\big)\Big\|\leq\frac{5\Delta_{T}}{\iota^{\prime}\iota_{T-1}}\end{split}$
which, together with (47), shows that
$\begin{split}\sup_{a\in\mathbb{R}}\Big|\Psi_{G_{c+c_{t-p}}}(\mu_{p})(h_{a,t-p})-\Psi_{G_{c+c_{t-p}}}\big(\hat{\mu}_{p-1}M\big)&(h_{a,t-p})\Big|\leq\frac{5\Delta_{T}}{\iota^{\prime}\iota_{T-1}}.\end{split}$ (50)
On the other hand, by Lemma 3 we have
and therefore, noting that $|x|G_{a}(x)\leq a^{-1/2}e^{-1/2}$ for all $x\in\mathbb{R}$ and all $a\in(0,\infty)$,
where the second inequality uses the fact that $\inf_{k\geq 1}\sigma^{2}_{k}=\sigma^{2}_{1}=\sigma^{2}/(1+c\sigma^{2})>0$ while $\inf_{k\geq 1}c_{k}\geq 0$ by Lemma 5.
Using the latter result and recalling that $c_{\star}=\rho^{2}/\sigma^{2}\geq\sup_{k\geq 1}c_{k}$, it follows that we have both
$\sup_{a\in\mathbb{R}}\Big|\Psi_{G_{c+c_{t-p}}}(\mu_{p})(h_{a,t-p})\Big|\leq\frac{|\rho_{t-p}|}{\sigma_{1}\,\sqrt{c}\,\mu_{p}(G_{c+c_{t-p}})}\leq\frac{|\rho_{t-p}|}{\sigma_{1}\,\sqrt{c}\,\mu_{p}(G_{c+c_{\star}})}$
and
$\sup_{a\in\mathbb{R}}\Big|\Psi_{G_{c+c_{t-p}}}\big(\hat{\mu}_{p-1}M\big)(h_{a,t-p})\Big|\leq\frac{|\rho_{t-p}|}{\sigma_{1}\,\sqrt{c}\,\hat{\mu}_{p-1}M(G_{c+c_{t-p}})}\leq\frac{|\rho_{t-p}|}{\sigma_{1}\,\sqrt{c}\,\hat{\mu}_{p-1}M(G_{c+c_{\star}})}$
where $\hat{\mu}_{p-1}M(G_{c+c_{\star}})\geq\iota^{\prime\prime}\iota_{T-1}$ with $\iota^{\prime\prime}=\eta_{1}(G_{c+c_{\star}})\inf_{x\in I}G(x)M(x,G_{c+c_{\star}})$. (This definition of $\iota^{\prime\prime}$ allows to cover both the case where $p=1$ and the case where $p>1$.) Consequently, if we define the constant $\bar{C}=\max\{5,2/(c\sigma_{1}^{2})^{1/2}\}$, then
$\begin{split}\sup_{a\in\mathbb{R}}\Big|\Psi_{G_{c+c_{t-p}}}(\mu_{p})(h_{a,t-p})-\Psi_{G_{c+c_{t-p}}}\big(\hat{\mu}_{p-1}M\big)(h_{a,t-p})\Big|\leq\frac{\bar{C}}{\iota^{\prime\prime}\iota_{T-1}}|\rho_{t-p}|.\end{split}$ (51)
By combining (46), (50) and (51), and recalling that $\bar{C}\geq 5$, we obtain that